OpenStudy (anonymous):

Which system of inequalities is shown in the graph? A) y > x – 7 and 3x – y > -1 B) y ≥ x – 7 and 3x – y < -1 C) y ≥ x – 7 and 3x – y ≥ -1 D) y ≤ x – 7 and 3x – y ≤ -1

OpenStudy (anonymous):

graph here

OpenStudy (trojanpoem):

Graph the inequalities.

OpenStudy (trojanpoem):

A) y = x - 7 , 3x - y = -1 y = 3x +1 when graphing inequalities you graph it as a simple line equation y = mx + b and shad the area represented.

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

ok

OpenStudy (trojanpoem):